# Lattice Boltzmann methods for soft flowing matter

### v1

@@ V.3 Dense emulsion under thermal flows
The Shan-Chen approach has been extensively used in applications, such as soft glassy systems Benzi et al. (2009, 2014); Sbragaglia et al. (2012) and coarsening dynamics Pelusi et al. (2019); Tiribocchi et al. (2019), where heat transfer is generally neglected. However, heat transfer can be relevant in a variety of contexts, ranging from polymeric mixtures Voit et al. (2005) and binary fluids Gonnella et al. (2010) to the convective motion of magma Stein and Spera (1992). In Refs.Pelusi et al. (2022, 2021), Pelusi and coworkers have extended the applicability of the Shan-Chen approach to the study of dense emulsions subject to thermal convection, where the temperature field is coupled with the momentum equation of the emulsion. More specifically, the dynamics of the scalar temperature $T$ is simulated by introducing an auxiliary set of distribution functions governed by a single relaxation time lattice Boltzmann equation
@@
A systematic investigation on the effect produced by varying droplet concentrations, from very dilute regimes (Newtonian emulsion) to dense ones, has been carried on in Ref.Pelusi et al. (2021). It has been shown, for example, that, at increasing concentration, the Nusselt number (measuring the heat transport efficiency) displays significant fluctuations around the mean value, in stark contrast with the dilute case where the fluctuations turn very weak and the transition to convection is accompanied by a steady flow. These results show that, alongside the near-contact LB method, the Shan-Chen approach can provide a further valuable platform to study the physics of confined dense emulsions where coalescence is inhibited. It would be of interest to extend these simulations to three dimensions where out-of-plane components of the fluid velocity are likely to play a non-trivial role.
Figure 17: (a) Initial setup of a dense emulsion arranged as a honeycomb structure. (b) Relaxed configuration, taken as a starting point for the convection simulation. In these configurations, the temperature does not affect the fluid flow. (c) Instantaneous configuration of a Rayleigh-Bénard convection in a cell once the emulsion is heated from below. Black arrows indicate the direction of the droplet and the color map represents the values of the temperature. Finally, ${\bf g}$ is the gravitational acceleration. The figure is reproduced with permission from Ref.Pelusi et al. (2022).
@@ V.4 Flowing hierarchical emulsions
An alternative numerical strategy to study the hydrodynamics of dense emulsions with hierarchical structure (a broad class including soft granular media as well) is based on the hybrid free energy LB method using a multiphase field approach (see section III.3.2). In this model, a set of scalar field $\phi_{n}({\bf x},t)$, $n=1,...,N$ describes the density of each droplet (being $N$ the total number of droplets) and the vector field ${\bf u}$ accounts for the global fluid velocity. The dynamics of the scalar fields is governed by a set of advection-diffusion equations of the form of Eq.(52), while the velocity field obeys the Navier-Stokes equation whose stress tensor includes forcing terms associated to each phase. Over the last few years this approach, often combined with LB methods, has been used to simulate systems such as deformable droplets in microchannels Foglino et al. (2017, 2018); Negro et al. (2023); Tiribocchi et al. (2020a) and cell monolayers displaying liquid crystal features Loewe et al. (2020); Mueller et al. (2019); Armengol-Collado et al. (2023). In this subsection, we discuss its use in the context of multiple emulsions flowing within a microfluidic channel, as presented in Ref.Tiribocchi et al. (2021c).
@@
Unlike the examples shown in subsections V.1 and V.2 focused on droplet production, here the interest is in the long-term behavior of these systems. The emulsion is initialized by defining a set of phase fields $\phi_{n}$ which relaxes towards equilibrium for a relatively short simulation time (see Fig.18, top row). At the boundaries, no-slip conditions are set for the fluid velocity and neutral wetting for the phase fields, although no wetting would essentially lead to similar results provided that the interaction between droplets and walls is negligible. Once mechanical equilibrium is attained, a constant pressure gradient is applied across the longitudinal direction of the channel. As time progresses, a large variety of non-equilibrium states is found, ranging from long-lived ones at low values of area fraction of the cores to short-lived ones at higher values. In the former regime, the cores are found to display a correlated planetary-like motion, in which they cease each other while remaining confined either in the upper or lower part of the emulsion for long periods of time. In the latter one, the increase of area fraction triggers collisions and multiple crossings between the two regions, finally leading to a chaotic-like dynamics.
@@
In the bottom row of Fig.18, we show a selection of these steady states. The first scenario is found at values of core area fraction lower than $0.35$ (corresponding to $N=1,2,3$), where the cores exhibit the periodic dynamics confined in a portion of the emulsion, while the second one, in which crossing events occur, appears at higher values (corresponding to $N=4,5,6$) up to approximately $0.5$. This complex behavior is decisively driven by the internal vorticity, consisting of two counter-rotating vortices within the external droplet. If the core area fraction is below $0.35$, the vortex structure remains essentially unaltered (such as the one shown in Fig.18m), whereas it turns chaotic for higher values, due to the non-trivial coupling of the fluid velocity with the interfaces of the cores.
@@
Figure 18: Top row: Equilibrium configurations of a double emulsion (a), a two-core (b), a three-core (c), a four-core (d), a five-core (e) and a six-core (f) emulsion. A set of scalar fields $\phi_{n}$ is associated to each droplet. Bottom row: (a)-(l) The figures show a list of non-equilibrium steady states of multiple emulsions under a pressure-driven flow. If the area fraction occupied by the cores is lower than $\sim 0.35$, the states are long-lived (l.l., corresponding to $N=1$, $N=2$, and $N=3$). The cores are either stuck at the front end or remain confined in the upper/lower part of the emulsion, often displaying a periodic motion along circular trajectories. If the area fraction is higher, the states turn to short-lived (s.l., corresponding to $N=4$, $N=5$, and $N=6$), and multiple crossings from one region towards the other one are observed. Here, Reynolds numbers range from $\sim 1$ to $\sim 3$ and Capillary numbers from $\sim 0.35$ to $\sim 0.85$. The dotted black arrow indicates the direction of motion of the emulsion, while the magenta one the direction of a crossing. In each snapshot, the steady states are classified in terms of an analogy with the occupation number formalism. (m) Typical double-vortex structure of the velocity field computed in the reference frame of a droplet subject to a pressure-driven flow. The color map ranges from 00 to $2$ and accounts for the values of the phase fields. Copyright granted under the https://creativecommons.org/licenses/by/4.0/ CC 4.0 International license. No changes were made from the original figure from Ref.Tiribocchi et al. (2021c).
@@
Production and dynamics of a double emulsion in a planar flow focusing can be also described using a ternary free-energy lattice Boltzmann Wöhrwag et al. (2018); Wang et al. (2020) where, unlike the hybrid version previously mentioned, the Navier-Stokes equation and the two equations of the order parameters (i.e. the densities of the inner and outer droplets) are solved by means of three different sets of distribution functions, essentially one for each hydrodynamic field. A selection of flow regimes obtained using this method for different values of Capillary numbers are shown in Fig.19, where a significant match with experiments is also reported. One can identify, for example, the two-step formation regime, (a1-b1), the one-step one (a2-b2), the decussate regime with one empty droplet (a3-b3), and the decussate regime with two empty droplets (a4-b4).
Figure 19: Experimental flow regimes reported in experimental Ref.Abate et al. (2011) (a1,a2) and simulation Azarmanesh et al. (2016) (a3-a4). Simulations from Ref.Wang et al. (2020) (b1-b4). Capillary numbers reported in Ref.Wang et al. (2020) are: (b1) $Ca_{i}=0.012$, $Ca_{m}=0.011$ and $Ca_{o}=0.035$; (b2) $Ca_{i}=0.018$, $Ca_{m}=0.011$ and $Ca_{o}=0.035$; (b3) $Ca_{i}=0.008$, $Ca_{m}=0.011$ and $Ca_{o}=0.035$; (b4) $Ca_{i}=0.008$, $Ca_{m}=0.011$ and $Ca_{o}=0.065$, where the subscripts $i$, $m$ and $o$ indicate inner, medium and outer fluid, respectively. Copyright granted under the https://creativecommons.org/licenses/by/4.0/ CC 4.0 International license. No changes were made from the original figure from Ref.Wang et al. (2020)
@@
The method discussed in this subsection has been also adopted to describe active fluids where, unlike the previous cases, spontaneous flows emerge at a mesoscale level as a result of the action of smaller units that self-propel Marchetti et al. (2013); Ramaswamy (2010). Typical examples are dense suspensions of bacteria dispersed in a fluid, protein networks in living cells, and artificial microswimmers. In the next section, we present a recent application of the lattice Boltzmann method to the dynamics of an active gel droplet migrating through a microfluidic constriction.
@@ V.5 Migration of active droplets through constrictions
The theoretical framework capturing the behavior of these systems at a continuum level (described in Section III.3.2) combined with the free-energy lattice Boltzmann has been particularly fruitful in simulating the dynamics of active fluid droplets in a highly diluted regime and under confinement. In Fig.20A-B (left panel), we show, for example, the structures of the polar and velocity field of a contractile (i.e. $\zeta<0$) fluid droplet, representing a minimal model of an actomyosin cell extract, in an unbounded medium. The droplet is initialized as a circular region where the polarization vector ${\bf p}$ is uniform and unidirectional. At increasing values of $\zeta$, two regimes can be distinguished: a non-motile one where a four-roll mill flow stretches the droplet longitudinally and a motile one where once $\zeta$ overcomes a critical value, the internal flow acquires a double-vortex pattern propelling the droplet forward. This structure breaks the inversion symmetry of the polar field, leading to an emergent elastic splay and motion along the direction of ${\bf p}$.
@@
A similar theoretical description can be used to study the dynamics of a contractile droplet through constrictions, a phenomenon of relevance in a number of biological processes such as cancer spreading and wound healing Davidson et al. (2015); Au et al. (2016); Cao et al. (2016); Brückner et al. (2019). In Fig.20a-v, we show the migration of a droplet of diameter $D$ through a pore modeled as two static and fluid-free phase fields (the pillars) placed at distance $h$ and glued to opposite flat walls. Although a crude approximation of a realistic constriction, this design allows for a simple implementation of mesoscale effects, such as repulsion and adhesion between the droplet and pillars. Adhesion, in particular, can be included by adding a contribution of the form $\sum_{n,m,n<m}\gamma_{n,m}\nabla\phi_{n}\nabla\phi_{m}$ to the free energy, where $\gamma_{n,m}$ is an adhesion constant between different phase fields.
@@
Importantly, adhesive forces are found to be crucial in enabling the crossing, especially through narrow interstices Tiribocchi et al. (2023d). Indeed, while for large pores (Fig.20a-f, $h/D\simeq 0.8$) the crossing generally occurs if the speed (which is controlled by the activity $\zeta$) is high enough, for smaller pores (Fig.20g-v, $h/D\simeq 0.5$ and $h/D\simeq 0.2$) the sole active stress is not sufficient to enable the process. In this case, the droplet would either get stuck at the pore’s entry or bounce back into the microchannel. Simulations show that, under these conditions, non-uniform adhesive forces (i.e. higher at the entry of the pore and lower at the exit) are necessary for a successful crossing, provided that inertia, surface tension, and adhesion are carefully balanced. Suitable dimensionless numbers assessing the importance of these effects are the adhesion number $A=\gamma_{L}/\gamma_{R}$, which is the ratio between adhesion forces at the entry and the exit of the pore, and the inertia-over-adhesion number $I_{A_{L,R}}=\rho v^{2}R^{2}/\gamma_{L,R}$. For $\lambda\simeq 0.5$, for example, one has $2\lesssim A\lesssim 10$, $I_{A_{R}}\simeq 0.25$ and $0.025\lesssim I_{A_{L}}\lesssim 0.06$. The values of the last two numbers, in particular, mean that inertial forces are weaker than adhesive ones at the entry (L) of the pore to enhance the connection between the droplet and pillars. A lower adhesiveness at the exit (R) allows for an easier detachment from the pore, finally favoring the crossing.
@@
Although this model remains distant from a living cell in many aspects, such as the lack of the nucleus, which represents a limiting factor in the crossing, yet these results may suggest that some features of cell motility can be reasonably captured by minimal mesoscopic models built upon a restricted number of dynamic parameters and hydrodynamic fields.
@@
Figure 20: Left panel, (A)-(B). In (A), the figure shows the initial configuration of the active droplet (left), the intermediate non-motile state (middle), and the motile state (right). Red arrows indicate the direction of the polar field, which turns from a uniform orientation at equilibrium to a splay deformation at the steady state. In (B), the figure shows the steady-state velocity fields for the intermediate configuration (left) and the motile one (right). The former consists of a four-roll mill flow, while the latter exhibits two counter-rotating vortices propelling the droplet. Here, red arrows indicate the direction of the velocity field. Right panel, (a)-(v). This panel shows the crossing of an active fluid droplet of diameter $D$ through a constriction of height $h$. In (a)-(f) $h/D\simeq 0.8$, in (g)-(n) $h/D\simeq 0.5$ and in (o)-(v) $h/D\simeq 0.2$. Decreasing the ratio $h/D$ yields larger shape deformations, some closely resembling experimentally observed structures, such as ampule-like (j,k,r,t) and hourglass (s) configurations. For $h/D\lesssim 0.5$, larger adhesive forces (pink layer in (g) and (o)) at the entrance of the pore and weaker ones (cyan layer) at the exit are crucial to favor the crossing. Here, $\gamma_{L}$ and $\gamma_{R}$ indicate the adhesion coefficient at the left and right sides of the pillars, large red arrows highlight splay deformations, and blue ones indicate bend distortions. The color map ranges from 00 to $2$ and accounts for the values of the phase fields. The left panel is reproduced with permission from PNAS Ref.Tjhung et al. (2012). Copyright granted under the https://creativecommons.org/licenses/by/4.0/ CC 4.0 International license. No changes were made from the original figure from Ref Tiribocchi et al. (2023d).

### v2

@@ V.3 Dense emulsion under thermal flows
The Shan-Chen approach has been extensively used in applications, such as soft glassy systems Benzi et al. (2009, 2014); Sbragaglia et al. (2012), or to simulate the coarsening dynamics Pelusi et al. (2019); Tiribocchi et al. (2019), where heat transfer is generally neglected. However, heat transfer can be relevant in a variety of contexts, ranging from polymeric mixtures Voit et al. (2005) and binary fluids Gonnella et al. (2010) to the convective motion of magma Stein and Spera (1992). In Refs.Pelusi et al. (2022, 2021), Pelusi and coworkers have extended the applicability of the Shan-Chen approach to the study of dense emulsions subject to thermal convection, where the temperature field is coupled with the momentum equation of the emulsion. More specifically, the dynamics of the scalar temperature $T$ is simulated by introducing an auxiliary set of distribution functions governed by a single relaxation time lattice Boltzmann equation
@@
A systematic investigation on the effect produced by varying droplet concentrations, from very dilute regimes (Newtonian emulsion) to dense ones, has been carried on in Ref.Pelusi et al. (2021). It has been shown, for example, that the Nusselt number (measuring the heat transport efficiency), time-averaged over statistical steady states, remains below that of a single phase system (i.e. when droplet concentration $\Phi_{0}$ tends to zero), is roughly constant for $\Phi_{0}\lesssim 0.2$ and progressively diminishes for higher values of $\Phi_{0}$. At increasing concentration, it displays significant fluctuations around the mean value, in stark contrast with the dilute case where the fluctuations turn very weak and the transition to convection is accompanied by a steady flow (see Fig.19d). Further simulations have characterized the rheological properties via a Couette cell, by confining the emulsion (now at constant temperature) between flat walls placed at distance $h$ and moving along opposite directions with velocity $u$, thus under a shear rate $\dot{\gamma}=2u/h$. The numerical results show that the measured relative viscosity $\eta_{r}=\eta_{eff}/\eta_{solv}$ (where $\eta_{eff}=d\Sigma/d\dot{\gamma}$ is the effective viscosity and $\eta_{solv}$ is the viscosity of the solvent) is in good agreement with values reported in literature for $\Phi_{0}\lesssim 0.12$, where $\eta_{r}$ follows a linear behavior proportional to $\Phi_{0}$ (for small droplet deformations) Einstein (1906); Taylor (1932). For larger values of $\Phi_{0}$, numerical data deviate from the linear behavior and are found to agree with results shown in Ref.Zinchenko (1984).
These results show that, alongside the near-contact LB method, the Shan-Chen approach can provide a further valuable platform to study the physics of confined dense emulsions where coalescence is inhibited. It would be of interest to extend these simulations to three dimensions where out-of-plane components of the fluid velocity are likely to play a non-trivial role.
Figure 19: (a) Initial setup of a dense emulsion arranged as a honeycomb structure. (b) Relaxed configuration, taken as a starting point for the convection simulation. In these configurations, the temperature does not affect the fluid flow. (c) Instantaneous configuration of a Rayleigh-Bénard convection in a cell once the emulsion is heated from below. Black arrows indicate the direction of the droplet and the color map represents the values of the temperature. Finally, ${\bf g}$ is the gravitational acceleration. (d) Time-average of the Nusselt number for different values of $\Phi_{0}$. The dotted line represent the Nusselt number for solid spherical particles (SP) in the very dilute regime, where $\Phi_{0}=0$. The inset shows time-averaged fluctuations of the Nusselt number $\Delta Nu=\langle(Nu(t)-\langle Nu\rangle_{t})^{2}\rangle_{t}^{1/2}$. The grey region indicates values of concentration where non-Newtonian effects emerge. (e) Comparison between the effective viscosity from simulations (red pentagons) and literature data from Taylor-Einstein Taylor (1932); Einstein (1906), Ghigliotti Ghigliotti et al. (2010) and Zinchenko Zinchenko (1984). The figure is reproduced with permission from Ref.Pelusi et al. (2022).
@@ V.4 Flowing hierarchical emulsions
An alternative numerical strategy to study the hydrodynamics of dense emulsions with hierarchical structure (a broad class including soft granular media as well) is based on the hybrid free energy LB method using a multiphase field approach (see section III.3.2). In this model, a set of scalar field $\phi_{n}({\bf x},t)$, $n=1,...,N$ describes the density of each droplet (being $N$ the total number of droplets) and the vector field ${\bf u}$ accounts for the global fluid velocity. The dynamics of the scalar fields is governed by a set of advection-diffusion equations of the form of Eq.(52), while the velocity field obeys the Navier-Stokes equation whose stress tensor includes forcing terms associated to each phase. Over the last few years the multi-phase field theory, often combined with LB methods, has been used to simulate systems such as deformable droplets in microchannels Foglino et al. (2017, 2018); Negro et al. (2023); Tiribocchi et al. (2020a) and cell monolayers displaying liquid crystal features Loewe et al. (2020); Mueller et al. (2019); Armengol-Collado et al. (2023). In this subsection, we discuss its use in the context of multiple emulsions flowing within a microfluidic channel, as presented in Ref.Tiribocchi et al. (2021c).
@@
Unlike the examples shown in sections V.1 and V.2 focused on droplet production, here the interest is in the long-term behavior of these systems. The emulsion is initialized by defining a set of phase fields $\phi_{n}$ which relaxes towards equilibrium for a relatively short simulation time (see Fig.20, top row). At the boundaries, no-slip conditions are set for the fluid velocity and neutral wetting for the phase fields, although no wetting would essentially lead to similar results provided that the interaction between droplets and walls is negligible. Once mechanical equilibrium is attained, a constant pressure gradient is applied across the longitudinal direction of the channel. As time progresses, a large variety of non-equilibrium states is found, ranging from long-lived ones at low values of area fraction of the cores to short-lived ones at higher values. In the former regime, the cores are found to display a correlated planetary-like motion, in which they cease each other while remaining confined either in the upper or lower part of the emulsion for long periods of time. In the latter one, the increase of area fraction triggers collisions and multiple crossings between the two regions, finally leading to a chaotic-like dynamics.
@@
In the bottom row of Fig.20, we show a selection of these steady states. The first scenario is found at values of core area fraction lower than $0.35$ (corresponding to $N=1,2,3$), where the cores exhibit the periodic dynamics confined in a portion of the emulsion, while the second one, in which crossing events occur, appears at higher values (corresponding to $N=4,5,6$) up to approximately $0.5$. This complex behavior is decisively driven by the internal vorticity, consisting of two counter-rotating vortices within the external droplet. If the core area fraction is below $0.35$, the vortex structure remains essentially unaltered (such as the one shown in Fig.20m), whereas it turns chaotic for higher values, due to the non-trivial coupling of the fluid velocity with the interfaces of the cores.Although a dedicated experiment on this complex dynamics is still missing, a behavior akin to the one observed for low core area fraction has been found in liquid crystal droplets moving inside microchannels Fernández-Nieves et al. (2007), where periodic director field changes are induced by the double-vortex flows inside the droplet. This effect, in turn, determines a cyclic motion of point defects which rotate in separate regions of the emulsion, closely resembling that of inner cores shown in Fig.20.
@@
Figure 20: Top row: Equilibrium configurations of a double emulsion (a), a two-core (b), a three-core (c), a four-core (d), a five-core (e) and a six-core (f) emulsion. Bottom row: (a)-(l) The figures show a list of non-equilibrium steady states of multiple emulsions under a pressure-driven flow. If the area fraction occupied by the cores is lower than $\sim 0.35$, the states are long-lived (l.l., corresponding to $N=1$, $N=2$, and $N=3$). The cores are either stuck at the front end or remain confined in the upper/lower part of the emulsion, often displaying a periodic motion along circular trajectories. If the area fraction is higher, the states turn to short-lived (s.l., corresponding to $N=4$, $N=5$, and $N=6$), and multiple crossings from one region towards the other one are observed. Here, Reynolds numbers range from $\sim 1$ to $\sim 3$ and Capillary numbers from $\sim 0.35$ to $\sim 0.85$. The dotted black arrow indicates the direction of motion of the emulsion, while the magenta one the direction of a crossing. In each snapshot, the steady states are classified in terms of an analogy with the occupation number formalism. (m) Typical double-vortex structure of the velocity field computed in the reference frame of a droplet subject to a pressure-driven flow. The color map ranges from 00 to $2$ and accounts for the values of the phase fields. Copyright granted under the CC 4.0 International license. No changes were made from the original figure from Ref.Tiribocchi et al. (2021c).
@@
Production and dynamics of a double emulsion in a planar flow focusing can be also described using a ternary free-energy LBWöhrwag et al. (2018); Wang et al. (2020) where, unlike the hybrid version previously mentioned, the Navier-Stokes equation and the two equations of the order parameters (i.e. the densities of the inner and outer droplets) are solved by means of three different sets of distribution functions, essentially one for each hydrodynamic field. A selection of flow regimes obtained using this method for different values of Capillary numbers are shown in Fig.21a-b, where a significant match with experiments is also reported. One can identify, for example, the two-step formation regime, (a1-b1), the one-step one (a2-b2), the decussate regime with one empty droplet (a3-b3), and the decussate regime with two empty droplets (a4-b4). A systematic study of the breakup modes leading to different flow regimes is shown in Fig.21c, where a three-dimensional phase diagram illustrates a combination of breakup modes of inner and middle fluids obtained by varying $Ca_{o}$, $Ca_{m}$ and $Ca_{i}$. Dripping-dripping, dripping-jetting and jetting-dripping modes, for example, are found for the two-step and the one-step formation shown in Fig.21a1-a2, experimentally observed in two-cross junctions and glass capillaries Abate et al. (2011); Kim et al. (2013). Further modes, such as dripping-threading, decussate and bidisperse and irregular-jetting are reported more frequently in glass capillary devices Nabavi et al. (2017); Kim et al. (2013); Shang et al. (2014).
Figure 21: (a-b) Flow regimes reported in experiments of Ref.Abate et al. (2011) (a1,a2) and simulation of Ref.Azarmanesh et al. (2016) (a3-a4) and Ref.Wang et al. (2020) (b1-b4). Capillary numbers reported in Ref.Wang et al. (2020) are: (b1) $Ca_{i}=0.012$, $Ca_{m}=0.011$ and $Ca_{o}=0.035$; (b2) $Ca_{i}=0.018$, $Ca_{m}=0.011$ and $Ca_{o}=0.035$; (b3) $Ca_{i}=0.008$, $Ca_{m}=0.011$ and $Ca_{o}=0.035$; (b4) $Ca_{i}=0.008$, $Ca_{m}=0.011$ and $Ca_{o}=0.065$, where the subscripts $i$, $m$ and $o$ indicate inner, medium and outer fluid, respectively. (c) Three dimensional phase diagram of different flow regimes observed by varying $Ca_{o}$, $Ca_{m}$ and $Ca_{i}$. The red solid line marks the transition from the dripping-mode breakup (left side) to jetting-mode breakup (right side) of the inner fluid. The blue dashed line separates a region (below the line) where the middle fluid breaks up in the dripping mode from a region (above the line) where the middle fluid breaks up in the jetting mode. Copyright granted under the CC 4.0 International license. The figures are adapted from Ref.Wang et al. (2020)
@@
The method discussed in this subsection has been also adopted to describe active fluids where, unlike the previous cases, spontaneous flows emerge at a mesoscale level as a result of the action of smaller units that self-propel Marchetti et al. (2013); Ramaswamy (2010). Typical examples are dense suspensions of bacteria dispersed in a fluid, protein networks in living cells, and artificial microswimmers. In the next section, we present a recent application of the LB method to the dynamics of an active gel droplet migrating through a microfluidic constriction.
@@ V.5 Migration of active droplets through constrictions
The theoretical framework capturing the behavior of these systems at a continuum level (described in Section III.3.2) combined with the free-energy LB method has been particularly fruitful in simulating the dynamics of active fluid droplets in a highly diluted regime and under confinement. In Fig.22A-B (left panel) we show, for example, the structures of the polar and velocity field of a contractile (i.e. $\zeta<0$) fluid droplet in an unbounded medium. The droplet is initialized as a circular region where the polarization vector ${\bf p}$ is uniform and unidirectional. At increasing values of $\zeta$, two regimes can be distinguished: a non-motile one where a four-roll mill flow stretches the droplet longitudinally and a motile one where, once $\zeta$ overcomes a critical value, the internal flow acquires a double-vortex pattern propelling the droplet forward. This structure breaks the inversion symmetry of the polar field, leading to an emergent elastic splay and motion along the direction of ${\bf p}$.
This 2D contractile droplet can be viewed as a simplified in silico version of an in-vitro cell extract, typically used in cell motility experiments. In this respect, the contractile-induced propulsion and the related flow structure could mimic the myosin-induced intracellular flow activating ketarocyte cells crawling on glass Yam et al. (2007). LB simulations also show that contractility can trigger autonomous motion of three-dimensional droplets Tjhung et al. (2012), a result that could be relevant for experiments of quasi-spherical tumor cells moving inside an elastic gel where the locomotion is shown to be caused mainly by myosin contractility Poincloux et al. (2011); Hawkins et al. (2011).
@@
A similar theoretical description can be used to study the dynamics of a contractile droplet through constrictions, a phenomenon of relevance in a number of biological processes such as cancer spreading and wound healing Davidson et al. (2015); Au et al. (2016); Cao et al. (2016); Brückner et al. (2019). In Fig.22a-v, we show the motion of a droplet of diameter $D$ through a pore, whose design draws inspiration from experiments of living cells (such as a mouse fibroblast) migrating within a microfluidic channel hosting constrictions Davidson et al. (2015, 2014). The pore is modeled by means of two static and fluid-free phase fields (the pillars) placed at distance $h$ and glued to opposite flat walls. Although a crude approximation of a realistic constriction, this design allows for a simple implementation of mesoscale effects, such as repulsion and adhesion between the droplet and pillars. Adhesion, in particular, can be included by adding a contribution of the form $\sum_{n,m,n<m}\gamma_{n,m}\nabla\phi_{n}\nabla\phi_{m}$ to the free energy, where $\gamma_{n,m}$ is an adhesion constant between different phase fields.
@@
Importantly, adhesive forces are found to be crucial in enabling the crossing, especially through narrow interstices Tiribocchi et al. (2023d). Indeed, while for large pores (Fig.22a-f, $h/D\simeq 0.8$) the crossing generally occurs if the speed (which is controlled by the activity $\zeta$) is high enough, for smaller pores (Fig.22g-v, $h/D\simeq 0.5$ and $h/D\simeq 0.2$) the sole active stress is not sufficient to enable the process. In this case, the droplet would either get stuck at the pore’s entry or bounce back into the microchannel. Simulations show that, under these conditions, non-uniform adhesive forces (i.e. higher at the entry of the pore and lower at the exit) are necessary for a successful crossing, provided that inertia, surface tension, and adhesion are carefully balanced. Suitable dimensionless numbers assessing the importance of these effects are the adhesion number $A=\gamma_{L}/\gamma_{R}$, which is the ratio between adhesion forces at the entry and the exit of the pore, and the inertia-over-adhesion number $I_{A_{L,R}}=\rho v^{2}R^{2}/\gamma_{L,R}$, where $v$ is the droplet velocity and $R$ is the droplet radius. For $\lambda\simeq 0.5$, for example, one has $2\lesssim A\lesssim 10$, $I_{A_{R}}\simeq 0.25$ and $0.025\lesssim I_{A_{L}}\lesssim 0.06$. The values of the last two numbers, in particular, mean that inertial forces are weaker than adhesive ones at the entry (L) of the pore to enhance the connection between the droplet and pillars. A lower adhesiveness at the exit (R) allows for an easier detachment from the pore, finally favoring the crossing.
@@
Although the match to experiments is mostly qualitative since this model remains distant from a living cell in many aspects (such as the lack of the nucleus, which represents a limiting factor in the crossing, and the absence of a membrane, replaced by a thin interface), these results may suggest that some features of cell motility can be reasonably captured by minimal mesoscopic models built upon a restricted number of dynamic parameters and hydrodynamic fields.
@@
Figure 22: Left panel, (A)-(B). In (A), the figure shows the initial configuration of the active droplet (left), the intermediate non-motile state (middle), and the motile state (right). Red arrows indicate the direction of the polar field, which turns from a uniform orientation at equilibrium to a splay deformation at the steady state. In (B), the figure shows the steady-state velocity fields for the intermediate configuration (left) and the motile one (right). The former consists of a four-roll mill flow, while the latter exhibits two counter-rotating vortices propelling the droplet. Here, red arrows indicate the direction of the velocity field. Right panel, (a)-(v). This panel shows the crossing of an active fluid droplet of diameter $D$ through a constriction of height $h$. In (a)-(f) $h/D\simeq 0.8$, in (g)-(n) $h/D\simeq 0.5$ and in (o)-(v) $h/D\simeq 0.2$. Decreasing the ratio $h/D$ yields larger shape deformations, some closely resembling experimentally observed structures, such as ampule-like (j,k,r,t) and hourglass (s) configurations. For $h/D\lesssim 0.5$, larger adhesive forces (pink layer in (g) and (o)) at the entrance of the pore and weaker ones (cyan layer) at the exit are crucial to favor the crossing. Here, $\gamma_{L}$ and $\gamma_{R}$ indicate the adhesion coefficient at the left and right sides of the pillars, large red arrows highlight splay deformations, and blue ones indicate bend distortions. The color map ranges from 00 to $2$ and accounts for the values of the phase fields. The left panel is reproduced with permission from Ref.Tjhung et al. (2012). Copyright granted under the CC 4.0 International license. No changes were made from the original figure from Ref Tiribocchi et al. (2023d).